A stationary observer heard an ambulance’s frequency 25 Hz lower than the 750 Hz at which it is emitted. If the air temperature is 20.0°c, how fast is the ambulance going?


1
Expert's answer
2021-01-26T13:15:41-0500

The Doppler's effect equation gives

"\Delta f=f_0\frac{v}{c}"

Hence,


"v=c\frac{\Delta f}{f_0}=\rm 340\: m/s\times\frac{25}{750}=11.3\: m/s"